A ball was kicked into the air from a balcony 20 feet above the ground, and the ball's height above the ground in feet t seconds after the ball was kicked is h(t) = 20 - 16t2 + 32t. What was the maximum height, in feet, of the ball above the ground after it was kicked?


Found 2 solutions by vjones93, Alan3354:
Answer by vjones93(1)
Answer by Alan3354(69443)

A ball was kicked into the air from a balcony 20 feet above the ground, and the ball's height above the ground in feet t seconds after the ball was kicked is h(t) = 20 - 16t2 + 32t. What was the maximum height, in feet, of the ball above the ground after it was kicked?
---------------
h(t) = (t) = - 16t2 + 32t + 20
Max ht is at the vertex of the parabola, at t = -b/2a
t = -32/-32 = 1
h(1) = -16 + 32 + 20
= 36 feet